OpenStudy (anonymous):

6. Simplify the expression. -3(b – 7)

OpenStudy (texaschic101):

do you know how to distribute the -3 through the parenthesis ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

you take the -3 and multiply it by everything in the parenthesis. what is -3 * b ? and what is -3 * - 7 ?

OpenStudy (texaschic101):

any idea ?

OpenStudy (anonymous):

i dont know what -3 * b is but ido know what -3*-7 is its -10

OpenStudy (texaschic101):

-3 * b = -3b -3 * - 7 = 21 (negative times a negative is a positive) so your answer is : -3b + 21 any questions ?

OpenStudy (anonymous):

yes thanks mate one second

OpenStudy (texaschic101):

we are not adding, we are multiplying, so -3 * -7 = 21

OpenStudy (anonymous):

i understand yup

OpenStudy (anonymous):

Solve the equation. Check your answer. m/20 = 5/4

OpenStudy (anonymous):

uhm im guessing its cross multiply

OpenStudy (texaschic101):

m/20 = 5/4 (this is a proportion) cross multiply (4)(m) = (20)(5) can you solve for m now ?

OpenStudy (anonymous):